OpenStudy (anonymous):

[6.03] What is the value of the x variable in the solution to the following system of equations? 4x − 3y = 3 5x − 4y = 3

OpenStudy (texaschic101):

4x - 3y = 3 -->(4)4x - 3y = 3 5x - 4y = 3 -->(-3)5x - 4y = 3 ------------ 16x - 12y = 12 (result of multiplying by 4) -15x + 12y = - 9 (result of multiplying by -3) --------------add x = 3 That is your answer but if you want to finish the problem just sub in 3 for x in either of the two equations, to find y. Then, if you want, you can check it by subbing in your known variables into either of the equations and if they equal, you have the correct answers :)